# A set $\mathcal{S}\subset\mathbb{N}$ is said to be a subset-sum-distinct or dissociated if all of its finite subsets have different sums. Alternately, an equivalent classification is if any equality of the form

### v1

@@ 1 Introduction
and $U(n):=\max\{k:\ B_{k}\leq n\}$. The asymptote above is an application of Stirling’s approximation. For a sharper two sided bound, one can use the so called Robbin’s estimate [26]. Also, let $c_{\ast}:=\frac{1}{2}\log_{2}\left(\frac{\pi}{2}\right)$.
@@
Fix a large $k$ and consider some $n\in\mathcal{I}_{k}:=[B_{k},,B_{k+1})$, and hence $U(n)=k$. Define
@@
and write $n=B_{k}r$ with $r\in[1,\rho_{k})$. Using $\rho_{k}=2(1+o(1))$, we may treat $r\in[1,2)$ asymptotically. Now, using Robbin’s estimate [26], we have
@@ 2 The Greedy Algorithm
This section is concerned with the greedy algorithm of generating dissociated sets. It starts with two given integers $a>0$ and $b>a$, and another integer $n$. It sets $\gamma_{1}=a$, $\gamma_{2}=b$ and in the $r$-th step it chooses $\gamma_{r}$ to be the smallest integer greater than $\gamma_{r-1}$ such that the sequence $\{\gamma_{k}\}_{k=1}^{r}$ is dissociated. The algorithm halts at the $n$-th step.
@@
Before going to the proof, we need some definitions and a few preliminary lemmas. Define For $A\subset\mathbb{N}$, we define
@@ 3 Generalization I
$\mathcal{S}(n)\,\leq\,\log_{2}n\,+\,\frac{1}{2}\log_{2}\log_{2}n\,+\,\log_{2}g\,-\,\log_{2}\left(\sqrt{\frac{2}{\pi}}\right)\,+\,o(1)$
@@
and hence $A_{n}\cup\{x\}$ is $\mathcal{D}[g]$ iff $r_{n}(s)+r_{n}(s-x)\leq g$ for all $s$ iff $x$ is not in the forbidden set
@@ 4 Generalization II
$\mathcal{S}(n)\,\leq\,\log_{k+1}n\,+\,\frac{1}{2}\log_{k+1}\log_{2}n\,+\,\frac{1}{2}\log_{k+1}\left(\frac{2\pi}{3}\,k(k+2)\right)\,+\,o(1)$
@@
By unimodality of $M_{m,k}$, we have (see [20])
@@
$\lim_{n\to\infty}\frac{\gamma_{n+1}^{(k)}}{\gamma_{n}^{(k)}}\to 2$

### v2

@@ 1 Introduction
and $U(n):=\max\{k:\ B_{k}\leq n\}$. The asymptote above is an application of Stirling’s approximation. For a sharper two sided bound, one can use the so called Robbins’ estimate [26]. Also, let $c_{\ast}:=\frac{1}{2}\log_{2}\left(\frac{\pi}{2}\right)$.
@@
Fix a large $k$ and consider some $n\in\mathcal{I}_{k}:=[B_{k},B_{k+1})$, and hence $U(n)=k$. Define
@@
and write $n=B_{k}r$ with $r\in[1,\rho_{k})$. Using $\rho_{k}=2(1+o(1))$, we may treat $r\in[1,2)$ asymptotically. Now, using Robbins’ estimate [26], we have
@@ 2 The Greedy Algorithm
This section is concerned with the greedy algorithm of generating dissociated sets. It starts with two given integers $a>0$ and $b>a$. It sets $\gamma_{1}=a$, $\gamma_{2}=b$ and in the $r$-th step it chooses $\gamma_{r}$ to be the smallest integer greater than $\gamma_{r-1}$ such that the sequence $\{\gamma_{k}\}_{k=1}^{r}$ is dissociated.
@@
Before going to the proof, we need some definitions and a few preliminary lemmas. Define For a finite $A\subset\mathbb{N}$, we define
@@ 3 Generalization I
$\left\lvert\mathcal{S}(n)\right\rvert\,\leq\,\log_{2}n\,+\,\frac{1}{2}\log_{2}\log_{2}n\,+\,\log_{2}g\,-\,\log_{2}\left(\sqrt{\frac{2}{\pi}}\right)\,+\,o(1)$
@@
and hence $\Gamma\big(\gamma_{n}\big)\cup\{x\}$ is $\mathcal{D}[g]$ iff $r_{n}(s)+r_{n}(s-x)\leq g$ for all $s$ iff $x$ is not in the forbidden set
@@ 4 Generalization II
$\left\lvert\mathcal{S}(n)\right\rvert\,\leq\,\log_{k+1}n\,+\,\frac{1}{2}\log_{k+1}\log_{2}n\,+\,\frac{1}{2}\log_{k+1}\left(\frac{2\pi}{3}\,k(k+2)\right)\,+\,o(1)$
@@
By unimodality of $M_{m,k}$ and Central Limit Theorem, we have (see [20]and [9])
@@
$\lim_{n\to\infty}\frac{\gamma_{n+1}^{(k)}}{\gamma_{n}^{(k)}}\to(k+1)$
@@ 5 Dissociated Sets Avoiding Geometric Progression
6 Acknowledgements
I would like to thank Prof. Ramachandran Balasubramanian, Prof. Greg Martin, Prof. Ilya Shkredov, Prof. Sándor Kiss, Satvik Saha and Sohom Gupta for the discussions I had with them about this topic.